# RhythmFormer: Extracting rPPG Signals Based on Hierarchical Temporal Periodic Transformer

### v1

@@ 4.3 Intra-dataset evaluation
PURE UBFC MMPD
Method Train Set MAE RMSE MAPE $\rho$ SNR MAE RMSE MAPE $\rho$ SNR MAE RMSE MAPE $\rho$ SNR
GREEN[45] - 10.09 23.85 10.28 0.34 -2.66 19.73 31.00 18.72 0.37 -11.18 21.68 27.69 24.39 -0.01 -14.34
ICA[34] - 4.77 16.07 4.47 0.72 5.24 16.00 25.65 15.35 0.44 -9.91 18.60 24.30 20.88 0.01 -13.84
CHROM[9] - 5.77 14.93 11.52 0.81 4.58 4.06 8.83 3.84 0.89 -2.96 13.66 18.76 16.00 0.08 -11.74
LGI[33] - 4.61 15.38 4.96 0.77 4.50 15.80 28.55 14.70 0.36 -8.15 17.08 23.32 18.98 0.04 -13.15
PBV[10] - 3.92 12.99 4.84 0.84 2.30 15.90 26.40 15.17 0.48 -9.16 17.95 23.58 20.18 0.09 -13.88
POS[48] - 3.67 11.82 7.25 0.88 6.87 4.08 7.72 3.93 0.92 -2.39 12.36 17.71 14.43 0.18 -11.53
DeepPhys[3] UBFC 5.54 18.51 5.32 0.66 4.40 - - - - - 17.50 25.00 19.27 0.06 -11.72
@@
PhysNet[51] UBFC 8.06 19.71 13.67 0.61 6.68 - - - - - 9.47 16.01 11.11 0.31 -8.15
PURE - - - - - 0.98 2.48 1.12 0.99 1.49 13.94 21.61 15.15 0.20 -9.94
@@
PURE - - - - - 1.30 2.87 1.50 0.99 1.49 13.94 21.61 15.15 0.20 -9.94
PhysFormer[56] UBFC 2.93 10.11 4.80 0.90 9.91 - - - - - 10.69 17.23 12.48 0.27 -8.84
PURE - - - - - 2.35 6.26 2.32 0.94 2.97 10.49 17.14 11.98 0.32 -8.27
@@
4.4 Cross-Dataset Evaluation
In addition, we followed the protocol outlined in [24] for cross-dataset evaluation. The models were trained on PURE and UBFC individually and subsequently tested on PURE, UBFC, and MMPD datasets. The training dataset was split, with the initial 80% used for training and the remaining 20% for validation. Comparative results against state-of-the-art end-to-end methods are presented in Table 3. The results highlight the superior capability of RhythmFormer in modeling domain-invariant rPPG features and its remarkable generalization across diverse datasets.
@@ 4.5 Ablation Study
We conducted ablation experiments on the MMPD dataset to assess the impact of different modules.
Table 4: Ablation of Fusion Stem.
Input Stem MAE RMSE MAPE $\rho$ SNR
@@
Table 5: Effectiveness of Fusion Stem.
@@
Impact of Fusion Stem. As illustrated in Table 4, through the comparison of ablation experiments with different input forms, we validated the effectiveness of the fusion stem. Compared to raw frame input and normalized frame difference input, the fusion stem significantly improves performance. Integrating the difference frame into the raw frame, enables frame-level awareness of BVP wave variations, effectively enhancing rPPG features. This further guides the self-attention mechanism of the transformer to reinforce its focus on rPPG representation.
@@
Furthermore, the fusion stem can be easily applied to other methods. We incorporated the fusion stem into existing methods. As demonstrated in Table 5, the application of the fusion stem to other methods led to significant performance improvements. This strongly attests to the effectiveness of the fusion stem, showcasing its capability to enhance rPPG features and improve signal-to-noise ratio across different methodologies.
Table 6: Ablation of Pre-Attention.
Pre-Attention MAE RMSE MAPE $\rho$ SNR
w/o 4.43 9.08 4.53 0.76 3.3
w. 3.07 6.81 3.24 0.86 5.46
Table 7: Ablation of Multi-Scale.
Mul-Scale Strategy n MAE RMSE MAPE $\rho$ SNR
w/o - 0 4.52 10.24 4.88 0.69 0.89
w/o - 1 4.90 11.07 4.99 0.65 1.73
w/o - 2 3.75 8.01 3.86 0.81 2.83
w/o - 3 4.26 8.85 4.42 0.76 3.57
w. concat 1,2,3 3.61 7.25 3.83 0.85 0.72
w. hierar. 3,2,1 5.10 10.73 5.21 0.64 0.96
w. hierar. 1,2,3 3.07 6.81 3.24 0.86 5.46
Impact of Pre-Attention. The main role of pre-attention is to reduce computational complexity and memory usage. Without pre-attention, a single batch’s memory usage surpasses the 48GB memory of a single RTX A6000, necessitating a reversion to coarse attention. As illustrated in Table 6, the results demonstrate that, with pre-attention, fine-grained rPPG information mining notably enhances modeling capability while mitigating noise interference.
@@
Method Param. (M) MACs (G)
DeepPhys[3] 7.504 120.00
PhysNet[51] 0.70 70.12
TS-CAN[22] 7.504 120.00
PhysFormer[56] 7.381 50.607
EfficientPhys[23] 7.439 60.687
RhythmFormer(ours) 3.251 38.494
@@
Impact of Multi-Scale. We compared the performance of the model under different fusion strategies and at different temporal scales, as shown in Table 7. The results in the first four rows indicate that, compared to a single scale, multi-scale modeling can more effectively capture the periodic rPPG features. The results in the $5_{th}$ row suggest that, unlike other periodic tasks(such as action counting [14]), hierarchical feature exploration is more effective than concatenation for extracting rPPG information that is of a single category and exhibits quasi-periodicity. The comparison between the last two rows shows that the multi-scale approach with an increasing temporal granularity is far superior to a decreasing temporal granularity. The findings suggest that reinforcing inter-frame rPPG features at a smaller time scale and then extracting periodic features at a larger time scale can effectively eliminate interference, enhancing the ability of the model to represent rPPG signals.
It is noteworthy that, within a single time scale, the use of a 4-frame granularity is significantly superior to others. Particularly, when the scale is set to 1-frame, i.e., globally attending to the temporal sequence, it incurs substantial memory and computational costs, yet the effect is suboptimal. This suggests that aggregating temporal information effectively enhances rPPG features and prevents noise interference.
4.6 Computational Cost
We also reported the parameters and multiply-accumulates (MACs) for all implemented methods, calculated under a video input size of 160×128×128 (T×H×W). The results are presented in Table 8. Our method achieves superior performance with fewer parameters and MACs, showcasing the potential for efficient mobile-level rPPG applications.
@@ 4.7 Visualization
As shown in Figure 3, we visualized the periodic sparse attention map. For the pre-attention stage, given a token (the $294_{th}$) located at the temporal trough and the spatially left face as a Query. The attention score map and key responses were obtained, where brighter regions indicate stronger attention. It provided the top-k (high-correlation) region for the refined attention stage to delve into rPPG features with finer granularity. For the refined attention stage, a sub-token (the $1179_{th}$) of the Query from the pre-attention was used as a Query in the top-k region. The attention score map and key responses were also obtained. As observed in Figure 3, the pre-attention score appropriately concentrates on skin regions with similar temporal phases, accurately expressing the periodic nature of rPPG. However, there are still many non-zero attention scores allocated to irrelevant areas. Refined attention focuses on the top-k regions, aiming to minimize attention on irrelevant areas as much as possible. This concentration of attention enables a more detailed extraction of rPPG features, reducing noise interference, and enhancing the model’s robustness.
@@
In response to the subtle and periodic nature of rPPG signals, we proposed a method that explicitly leverages periodicity to design the periodic sparse attention mechanism, and combines multi-scale temporal learning to unearth the quasi-periodic patterns in weak signals. This method achieved robust performance in rPPG measurement, surpassing previous state-of-the-art approaches across various datasets, and demonstrated outstanding generalization in cross-dataset evaluation. We envision RhythmFormer to serve as a new baseline in the fully supervised rPPG domain. Moreover, RhythmFormer proves highly effective in heart rate assessment tasks, and future exploration will delve into its feasibility for other tasks involving the extraction of periodic signals.
Acknowledgments This work was supported in part by the National Key Research and Development Program of China (2022ZD0117900), the National Natural Science Foundation of China (62206015, 62227801, U19B2032, U20B2062), the Guangdong Basic and Applied Basic Research Foundation (2021A1515110249), and the Fundamental Research Funds for the Central Universities (FRF-TP-22-043A1).
Supplementary Material
A Details of Data Augmentation
As shown in figure a, the heart rate (HR) distribution among the datasets is highly imbalanced, and both the PURE and UBFC datasets have limited amounts of data. Due to this, we employed a random sampling data augmentation strategy during training on PURE and UBFC datasets. For cases where the heart rate, calculated from the ground truth BVP wave, exceeded 90 beats per minute, random downsampling was applied; for rates below 75 beats per minute, random upsampling was performed. Additionally, random horizontal flips were applied to all videos.
Figure a: Dataset Heart Rate Distribution.
B Ablation study of HR Hybrid Loss
We conducted the ablation study on the MMPD dataset to assess the impact of the HR Hybrid Loss. The incorporation of the proposed HR constraint into the two commonly used losses from previous works. One constrained only in the time domain $\mathcal{L}_{MSE}$ and another employs negative Pearson loss in the time domain and cross-entropy loss in the frequency domain ${\mathcal{L}_{Hybrid}\ =\ \alpha\ \ast\ \mathcal{L}}_{Pearson}+\beta\ \ast\ \mathcal{L}_{CE}$. As shown in Table a, the results demonstrate that the HR Hybrid Loss achieves better performance by mitigating the gap between BVP loss and HR metrics, guiding the network to select a more optimal model.
We also computed the Pearson correlation coefficient between the loss and the metrics on the validation set, resulting in 0.81 for the proposed loss and 0.71 for the baseline. Additionally, as depicted in Figure b, the trend of the proposed loss during the later stages of training notably resembled that of the metrics, both reaching their nadir at the $27_{th}$ epoch.
C Visualization of the results
As shown in figure f, we present two examples of HR estimation using the proposed method, comparing the predicted BVP signals and the ground truth BVP signals in both the temporal and frequency domains. The predicted signals exhibit a strong alignment with the ground truth signals in terms of trends, yet they lack detailed high-frequency information.
To further assess the consistency between the predicted HR and the ground truth HR, the Bland-Altman plots for all the face videos across three experiments are respectively depicted in figure c,d,e. From the three figures, it can be observed that the scatter plot aligns overall with the y = x line, the points in the difference plot mostly fall within the confidence interval, but there are still some significant outliers with considerable deviations.
Table a: Ablation of HR Hybrid Loss
UBFC - PURE MMPD
loss MAE RMSE $\rho$ MAE RMSE $\rho$
$\mathcal{L}_{MSE}$ 2.47 8.80 0.93 3.15 7.10 0.85
$\mathcal{L}_{MSE}+\mathcal{L}_{HR}$ 0.99 3.01 0.99 3.15 7.10 0.85
$\mathcal{L}_{Hybrid}$ 1.84 7.82 0.94 3.13 6.98 0.85
$\mathcal{L}_{Hybrid}+\mathcal{L}_{HR}$ 0.97 3.36 0.99 3.07 6.81 0.86
Figure b: Visualization of the training process. The metrics are calculated on the validation set.
Figure c: The Bland-Altman Plots of Intra-dataset on MMPD.
Figure d: The Bland-Altman Plots of training on UBFC and testing on PURE.
Figure e: The Bland-Altman Plots of training on PURE and testing on UBFC.
(a) A example of Intra-dataset on PURE.
(b) A example of Intra-dataset on MMPD.
Figure f: Two examples of HR estimation using the proposed method. The frequency domain range corresponds to the range of heart rates.

### v3

@@ 4.5 Cross-dataset Evaluation
PURE UBFC-rPPG MMPD
[…]
GREEN [1] - 10.09 23.85 10.28 0.34 -2.66 19.73 31.00 18.72 0.37 -11.18 21.73 27.72 24.44 -0.02 -14.34
ICA [53] - 4.77 16.07 4.47 0.72 5.24 16.00 25.65 15.35 0.44 -9.91 18.57 24.28 20.85 0.00 -13.84
CHROM [14] - 5.77 14.93 11.52 0.81 4.58 4.06 8.83 3.84 0.89 -2.96 13.63 18.75 15.96 0.08 -11.74
LGI [54] - 4.61 15.38 4.96 0.77 4.50 15.80 28.55 14.70 0.36 -8.15 17.02 23.28 18.92 0.04 -13.15
PBV [55] - 3.92 12.99 4.84 0.84 2.30 15.90 26.40 15.17 0.48 -9.16 17.88 23.53 20.11 0.09 -13.88
POS [8] - 3.67 11.82 7.25 0.88 6.87 4.08 7.72 3.93 0.92 -2.39 12.34 17.70 14.43 0.17 -11.53
Face2PPGomit [9] - 4.66 15.82 4.97 8.76 4.37 16.99 29.54 15.91 0.34 -7.29 17.02 23.23 18.89 0.04 -12.77
DeepPhys [2] UBFC 5.54 18.51 5.32 0.66 4.40 - - - - - 17.50 25.00 19.27 0.05 -11.72
@@
PhysNet [19] UBFC 8.06 19.71 13.67 0.61 6.68 - - - - - 10.24 16.54 12.46 0.29 -8.15
PURE - - - - - 0.98 2.48 1.12 0.99 1.09 13.22 19.61 14.73 0.23 -10.59
@@
PURE - - - - - 1.30 2.87 1.50 0.99 1.49 13.94 21.61 15.14 0.20 -9.94
PhysFormer [11] UBFC 12.92 24.36 23.92 0.47 2.16 - - - - - 12.10 17.79 15.41 0.17 -10.53
PURE - - - - - 1.44 3.77 1.66 0.98 0.18 14.57 20.71 16.73 0.15 -12.15
@@
Table 3: Cross-dataset results.
The scatter plots and Bland-Altman plots in Figure 9 further demonstrate the strong correlation between the predictions and the ground truth. It is also observed that the outliers are predominantly located in regions corresponding to the values with the least amount of data in the training set.
@@ 4.6 Ablation Study
We conducted ablation experiments on the MMPD dataset to evaluate the impact of different modules.
Impact of Fusion Stem. As illustrated in Table 4, through the ablation study with different input forms, we validated the effectiveness of the fusion stem. Compared to raw frame input and normalized frame difference input, the fusion stem significantly improves performance. Integrating the frame differences into the raw frames enables frame-level representation to capture BVP wave variations, effectively enhancing rPPG features. This further guides the self-attention mechanism of the transformer to reinforce its focus on rPPG representation.
[…]
@@
Table 4: Ablation of fusion stem.
@@
Table 5: Effectiveness of fusion stem.
@@
Furthermore, the fusion stem can be easily applied to other methods. We incorporated the fusion stem into existing methods. As demonstrated in Table 5 and Table 2, the application of the fusion stem to other methods achieved consistent performance improvements. This strongly attests to the effectiveness of the fusion stem, showcasing its capability to enhance rPPG features and improve signal-to-noise ratio across different methods.
We conducted a detailed ablation study on the fusion stem. Table 6 reports the ablation results based on the number of adjacent frames, while another part of Table 7 reports the ablation results based on the step of adjacent frames. The best performance was achieved when using two surrounding frames with a step of 1. Steps of 0.5 and 1.5 involved cubic spline interpolation, which may explain the slight performance drop. The ablation results from the frame step also indicate that small variations in fps do not significantly affect the model’s performance, which is consistent with the results from the COHFACE dataset in Table 12.
[…]
0 4.56 9.44 4.60 0.73 0.68
1 4.19 8.90 4.29 0.77 1.94
2 3.07 6.81 3.24 0.86 5.46
3 3.74 8.24 3.86 0.79 4.26
Table 6: Ablation of frame number.
[…]
0.5 4.24 9.47 4.40 0.73 3.47
1.0 3.07 6.81 3.24 0.86 5.46
1.5 3.45 8.19 3.47 0.80 4.45
2.0 3.15 6.51 4.28 0.79 4.60
Table 7: Ablation of frame step.
[…]
$0:10$ 4.39 9.10 4.55 0.76 4.08
$1:9$ 3.45 7.77 3.56 0.82 5.16
$3:7$ 3.85 8.13 4.02 0.80 4.41
$5:5$ 3.07 6.81 3.24 0.86 5.46
$7:3$ 3.36 7.99 3.46 0.81 4.42
$9:1$ 3.27 7.12 3.43 0.85 4.71
$10:0$ 4.56 9.44 4.60 0.73 0.68
Table 8: Ablation of fusion ratio
Table 8 presents the ablation study of the fusion ratio when integrating the raw frames with the frame differences. The optimal results are achieved with a 5:5 fusion ratio. Furthermore, the first and last rows indicate that the fusion of raw frames and frame differences significantly enhances the rPPG feature representation, thereby improving the recovery quality of the BVP waveform.
Impact of Periodic Sparse Attention. As illustrated in Table 9, the result demonstrates that, compared to the vanilla attention mechanism, the periodic sparse attention mechanism achieves a finer-grained context relevance computation through pre-attention, significantly enhancing the modeling capability while reducing noise interference.
[…]
Vanilla Attention 4.43 9.08 4.53 0.76 3.30
Periodic Sparse Attention 3.07 6.81 3.24 0.86 5.46
Table 9: Ablation of periodic sparse attention.
[…]
1 - 0 4.52 10.24 4.88 0.69 0.89
1 - 1 4.71 10.63 4.97 0.67 0.93
1 - 2 3.75 8.01 3.86 0.81 2.83
1 - 3 4.26 8.85 4.42 0.76 3.57
2 hierarchical 1,2 4.07 8.86 4.19 0.77 4.07
3 concat 1,2,3 3.61 7.25 3.83 0.85 0.72
3 hierarchical 3,2,1 5.10 10.73 5.21 0.64 0.96
3 hierarchical 1,2,3 3.07 6.81 3.24 0.86 5.46
Table 10: Ablation of multi-temporal scale.
Impact of Multi-temporal Scale. We compared the performance of the model under different strategies and at different temporal scales, as shown in Table 10. Here, n represents the downsampling factor of the TPT block. The results in the first four rows indicate that, compared to a single temporal scale, multi-scale modeling can more effectively capture the periodic rPPG features.
Since the frame rate is 30 fps, when the number of TPT blocks exceeds 3, the downsampling factor becomes too large to capture a single heartbeat at relatively high heart rates. Therefore, the ablation study was conducted only for TPT block values up to 3. The results from the $2_{nd}$, $5_{th}$, and $8_{th}$ rows demonstrate a continuous improvement in performance as the scale increases. This may be due to the varying effects of noise at different scales, and the multi-scale approach can reduce the weight of noise at a single scale. We have plotted the Bland-Altman plots for these three rows in Figure 10 and also provided the specific waveforms for the same outlier sample. It can be observed that the distributions of the three Bland-Altman plots are quite similar, but the multi-scale approach effectively mitigates some outliers by reducing noise, leading to a smaller 95% confidence interval. From the waveforms, it is clear that the multi-scale method enhances important features while suppressing noise.
(a) Table 10, $2_{nd}$ row
(b) Table 10, $5_{th}$ row
(c) Table 10, $8_{th}$ row
Figure 10: Visualization of the ablation study on multi-temporal scale.
The results in the $6_{th}$ row suggest that, unlike other periodic tasks (such as action counting [34]), hierarchical feature extraction is more effective than concatenation for extracting rPPG signals, which belong to a single category and exhibit quasi-periodicity. Meanwhile, the comparison between the last two rows shows that the multi-scale approach with increasing temporal granularity significantly outperforms the approach with decreasing temporal granularity. This may be because noise typically manifests as high-frequency signals, and fine-grained features are more susceptible to noise interference. The abstract process from fine to coarse naturally ignores detailed noise and focuses on semantic information. The pyramid structure uses higher-level features (global context) to guide the optimization of lower-level features, enhancing important features and suppressing noise. As shown in the $5_{th}$ row, although the heart rate derived from the pulse wave is relatively inaccurate, the SNR of the reconstructed BVP compared to the ground truth BVP is relatively high, which supports this point.
It is noteworthy that, for a single temporal scale, a granularity of 4-frame (n=2) is significantly superior to others. Particularly, when the scale is set to 1-frame (n=0), i.e., globally attending to the temporal sequence, it incurs substantial memory and computational costs, yet the effect is suboptimal. This suggests that aggregating temporal information effectively enhances rPPG features and prevents noise interference.
Impact of Input Length. As shown in Table 11, an input length of 160 frames achieves the optimal results. This may be because face detection is performed only on the first frame. If the input length is too long, subsequent frames may lose the face due to motion in certain scenarios. In contrast, shorter input lengths may not effectively capture inter-frame rPPG signals.
[…]
80 3.53 7.77 3.64 0.82 2.74
160 3.07 6.81 3.24 0.86 5.46
240 3.34 7.12 3.57 0.85 6.03
320 3.86 7.81 4.09 0.83 4.15
Table 11: Ablation of input length.
Impact of FPS on the COHFACE. The original frame rate of the COHFACE dataset is 20 fps. Here, we conducted a comparative experiment by upsampling to 30 fps. As shown in Table 12, after upsampling, the MAE, RMSE, and SNR deteriorated, while the MAPE and $\rho$ improved. This indicates that upsampling allows the model to better capture the overall variation patterns but also introduces noise, reducing the model’s robustness.
[…]
20 1.17 3.36 1.84 0.97 9.28
30 1.54 3.60 1.51 0.98 5.61
Table 12: Impact of FPS on the COHFACE.
Impact of Loss Term and Bias Term. We conducted ablation experiments on each loss term and bias term, and the results indicate that the temporal loss is the core component, while the frequency-domain loss and HR bias also play a role.
[…]
3.56 8.17 3.61 0.80 4.19
13.32 16.54 14.15 0.38 -12.7
3.13 6.98 3.39 0.85 5.56
3.07 6.81 3.24 0.86 5.46
Table 13: Ablation of loss term and bias term.
4.7 Computational Cost
We conducted a 30-second inference at a resolution of 128$\times$128 for all implemented methods, reporting the parameter count, average multiply-accumulates (MACs) per frame, average throughput per frame, and average peak GPU memory per frame. The notation “(10s)” refers to the average results for inferring 10 seconds. We found that when inferring 10 seconds, the average computational cost per frame of CNN-based methods remained unchanged, while the throughput and GPU memory usage of transformer-based methods varied. GPU memory usage significantly increased due to the pairwise attention computation for tokens in transformer-based methods. The throughput analysis is provided in Section 4.9.
Methods Parameters (M) MACs (M) Throughput(Kfps) GPU Memory(M)
DeepPhys [2] 1.98 744.45 5.65 37.28
PhysNet [19] 0.77 438.24 11.80 11.43
TS-CAN [17] 1.98 744.45 5.21 38.91
EfficientPhys [23] 1.91 373.72 8.42 26.68
PhysFormer [11] 7.38 316.29 8.50 28.63
PhysFormer (10s) 7.38 316.29 10.27 22.17
Ours 3.25 240.55 6.30 29.06
Ours (10s) 3.25 240.55 9.44 19.07
@@ 4.8 Visualization
We visualized the periodic sparse attention map in Figure 11. For the pre-attention stage, given a token (the $294_{th}$) located at the temporal trough and the left face as a Query. The attention score map and key responses were obtained, where brighter regions indicate stronger attention. It provided the top-k (high-correlation) region for the refined attention stage to delve into rPPG features with finer granularity. For the refined attention stage, a sub-token (the $1179_{th}$) of the Query from the pre-attention was used as a Query in the top-k region. The attention score map and key responses were also obtained. As observed in Figure 11, the pre-attention score appropriately concentrates on skin regions with similar temporal phases, accurately expressing the periodic nature of rPPG. However, there are still many non-zero attention scores allocated to irrelevant areas. Refined attention focuses on the top-k regions, aiming to minimize attention on irrelevant areas as much as possible. This concentration of attention enables a more detailed extraction of rPPG features, reducing noise interference, and enhancing the model’s robustness.
4.9 Limitation
We leverage the temporal attention sparsity induced by periodicity and introduce pre-attention to learn the periodic patterns. Although pre-attention operates at a coarse region level and does not generate significant computation, it inevitably leads to additional GPU kernel launches and memory transactions. From Table 14, RhythmFormer has lower computational complexity compared to Physformer, but its throughput is lower, likely due to this effect. Nevertheless, this issue could be alleviated through engineering efforts, such as GPU kernel fusion.
@@ 5 Conclusion
In response to the subtle and periodic nature of rPPG signals, we propose a method that explicitly leverages periodicity to design a periodic sparse attention mechanism. Combined with the proposed plug-and-play fusion stem to guide attention, this method effectively recognizes quasi-periodic patterns in weak rPPG signals, achieving state-of-the-art performance in extensive experiments. We envision RhythmFormer as a baseline for the rPPG community, supporting diverse methods and applications in the rPPG domain. In the future, it is potential to explore multimodal periodic sparse attention, as it may align periodic patterns across various modalities, such as RGB, radar, and radio frequency. Additionally, periodic sparse attention has proven highly effective in the rPPG task, and we will investigate its feasibility for broader fine-grained or periodic video understanding tasks.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (62206015, 62227801, 62376024, 62402033), the National Science and Technology Major Project (2022ZD0117901), the Fundamental Research Funds for the Central Universities (FRF-TP-22-043A1), and the Young Scientist Program of The National New Energy Vehicle Technology Innovation Center (Xiamen Branch).